OpenStudy (iloveeenikooo):

medal Below are the data collected from two random samples of 500 American adults on the number of hours they work per day (rounded to the nearest hour): Number of hours of work per day 6 7 8 9 10 Sample A: Number of adults 60 90 145 150 55 Sample B: Number of adults 70 80 140 145 65 Ryan concludes that adults spend a mean of 8 hours working each day. Malia thinks the mean is 9 hours. Who is correct—Ryan or Malia? Explain your answer in two or three sentences. Make sure to use facts to support your answer.

OpenStudy (iloveeenikooo):

@Qwertty123 @Wolfboy @TheSmartOne @inkyvoyd @Opcode @princesssleelee @aaronq @adrynicoleb @Aveline @Awolflover1 @sleepyjess @sweetburger @jabez177 @Jack_Prism @KendrickLamar2014 @Vocaloid @BruhObamaCare @marcelie

OpenStudy (bruhobamacare):

What is 6(60) + 7(90) + 8(145) + 9(150) + 10(55)

OpenStudy (bruhobamacare):

6(60) + 7(90) + 8(145) + 9(150) + 10(55) = 4050/500 , so the mean of Sample A is 4050/500

OpenStudy (bruhobamacare):

And 6(70)+7(80)+8(140)+9(145)+10(65) = 4055/500 , so the mean of Sample B is 4055/500

OpenStudy (bruhobamacare):

And the mean of both means is 8.1

OpenStudy (iloveeenikooo):

thanks! @BruhObamaCare how should i put that into a few short sentences?

OpenStudy (iloveeenikooo):

Please just get the first sentence started at least lol I don't know what to say ):)

OpenStudy (iloveeenikooo):

@BruhObamaCare

OpenStudy (bruhobamacare):

Mean of Sample A is 4050/500 Mean of Sample B is 4055/500 Mean of both means is 8.1

OpenStudy (iloveeenikooo):

thanks so much!

OpenStudy (bruhobamacare):

No problem :D Glad I could help